The diameter of the field of view of a microscope at 40X total magnification is 4.5 mm

Suppose you were viewing an object using a total magnification of 250X.

a) What is the field of view?

b) If your object takes up 1/4 of the field at 250X, what is the approximate size of your object?
Answer #1

Magnification1 × diameter of field of view1 = magnification2 × diameter of field of view2

4.5 × 40x = diameter2 × 250x

Diameter2 = 0.72 mm

A. Area of field of view 2 = πr^2 = 22/7 × (0.72/2)^2 = 0.4073 mm^2

B. Size of object = 1/4 × 0.4073 = 0.1018 mm^2

Please rate high.